aleyshahp:

Ada is showing her work in simplifying (5.9 – 1.8) – 6.2 + 4.1. In which step did Ada make an error? Step 1: (5.9 – 1.8) – 6.2 + 4.1 Step 2: 5.9 + (–1.8 – 6.2) + 4.1 (associative property) Step 3: 5.9 – 8 + 4.1 Step 4: 5.9 + 4.1 – 8 (associative property) Step 5: 10 – 8 = 2 Step 2; she wrote associative instead of commutative Step 2; she wrote associative instead of distributive Step 4; she wrote associative instead of commutative Step 4; she wrote associative instead of distributive

itsme:

she used the associative property twice that may have sum thing to do with it @supie

supie:

The way you would solve it is \[5.9−1.8−6.2+4.1=\]\[4.1−6.2+4.1=\]\[−2.1+4.1=\]\[2\]

aleyshahp:

thx

supie:

np even though that was a month ago XD